OpenStudy (anonymous):

Find the equation of the normal line (in point slope form) to f(x)= xcos(x)+x x=2pi Please give a full explanation!! Thanks!

OpenStudy (loser66):

your idea?

OpenStudy (anonymous):

I tried using the definition of a derivative but that got too complicated, and then Is searched a different method online and I found that the slope is 2 but 2cos2+2 seemed weird

OpenStudy (loser66):

f' (x ) = ?

OpenStudy (anonymous):

f'(x)= -xsinx+cosx+1

OpenStudy (loser66):

so, f' ( 2pi) = ?

OpenStudy (anonymous):

well -2pi(sin2pi)=0, and cos(2pi)+1=2, so 2?

OpenStudy (loser66):

yup

OpenStudy (loser66):

one more thing, when x = 2pi, f(2pi) = ?

OpenStudy (anonymous):

but isn't that just the slope?

OpenStudy (loser66):

not yet, be patient,

OpenStudy (anonymous):

I had trouble finding the (x,y) to put into point slope form

OpenStudy (loser66):

answer me, x = 2pi, f( 2pi) = ?

OpenStudy (anonymous):

so xcosx+x --> 2pi(cos2pi) + 2pi, so 4pi?

OpenStudy (loser66):

so, you have a point, that is (2pi, 4pi) and you have a slope of TANGENT line is 2, And you need a NORMAL line which is perpendicular to Tangent line, right? and the slope of that NORMAL line is -1/2, got me?

OpenStudy (anonymous):

just a question here, normal line=tangent?

OpenStudy (loser66):

perpendicular to tangent,

OpenStudy (anonymous):

Yes I got you!

OpenStudy (loser66):

so, DAT SIT

OpenStudy (anonymous):

so the normal line would be y-4pi=2(x-2pi)

OpenStudy (anonymous):

I meant 1/2 for the slope lol

OpenStudy (loser66):

-1/2, not 1/2

OpenStudy (anonymous):

thank you so much!!!

OpenStudy (loser66):

hopefully I don't make any mistake, If there is, someone points it out , please